Principal of $1700 is invested at 8.2% simple interest, paid annually, for 7 years
Answer by JulietG(1812)

In the first year, 1700 * 1.082 = 1839.4
In the second year, 1839.4 * 1.082 = 1990.23
In the third year, 1990.23 * 1.082 = 2153.43
and so on.
There's a simple interest calculator at http://www.webmath.com/simpinterest.html